OpenStudy (theedwardsfamily):

Which ratios are proportional to 27/30 ? Choose all answers that are correct. A. 3/10 B. 18/20 C. 54/60 D. 81/90 @Legends @Holly00d1248

OpenStudy (theedwardsfamily):

@Godlovesme

OpenStudy (theedwardsfamily):

@samanthagreer @sammixboo

OpenStudy (anonymous):

First reduce 27/30 to lowest terms. Are there any numbers that divide evenly into both 27 and 30?

OpenStudy (theedwardsfamily):

3 and 5?

OpenStudy (holly00d1248):

no 5 doesnt only 3

OpenStudy (anonymous):

Does 5 divide evenly into 27?

OpenStudy (theedwardsfamily):

no

OpenStudy (legends):

does 3 go into 27?

OpenStudy (theedwardsfamily):

yea

OpenStudy (holly00d1248):

I'm pretty sure it B @TheEdwardsFamily

OpenStudy (anonymous):

Correct. So we can rewrite as follows;\[\frac{ 27 }{ 30 }=\frac{ 3\times 9 }{ 3\times 10 } = \frac{ 9 }{ 10 }\] Understand so far?

OpenStudy (theedwardsfamily):

could there be anothr answr?

OpenStudy (holly00d1248):

do you mean two answers

OpenStudy (legends):

ok so 3 goes into 27 does 10 go into 30

OpenStudy (theedwardsfamily):

yes

OpenStudy (anonymous):

Yes, There is more than one correct answer.

OpenStudy (theedwardsfamily):

which would tht be?

OpenStudy (holly00d1248):

I told you wrong @TheEdwardsFamily one answer is D not B

OpenStudy (anonymous):

I you understood that 27/30 in lowest terms is 9/10, then can you multiply both numerator and denominator by the same number to get any of the choices? Take choice 1 for example. Can you multiply both numerator and denominator of 9/10 by the same number to get 3/10?

OpenStudy (theedwardsfamily):

ok. is it D and wht? i confused

OpenStudy (holly00d1248):

I only see one answer @ospreytriple

OpenStudy (theedwardsfamily):

but is has to be more than 1

OpenStudy (anonymous):

Does it SAY that it NEEDS to be more than one? Because sometimes they trick you with the multiple choice thing

OpenStudy (theedwardsfamily):

yes.

OpenStudy (holly00d1248):

Why???

OpenStudy (anonymous):

Consider choice 2. Can you multiply both the numerator and denominator of 9/10 by the same number to get 18/20?

OpenStudy (holly00d1248):

Well than im stumped because i only see one answer

OpenStudy (anonymous):

I see two answers

OpenStudy (anonymous):

Let me give you guys a hint 27 * 3 =

OpenStudy (holly00d1248):

O i see two answers now your ight @HailKK and @ospreytriple

OpenStudy (theedwardsfamily):

81

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

There are three correct answers. Sorry folks. I'm outta here. Too many people not helping, just confusing the poor student. You may message me if you want additional help.

OpenStudy (anonymous):

I see 3 answers

OpenStudy (anonymous):

FINE THERE ARE THREE

OpenStudy (theedwardsfamily):

I have to complet this before 6:00 pm

OpenStudy (anonymous):

Its BCD

OpenStudy (theedwardsfamily):

r u sure??

OpenStudy (holly00d1248):

Please explain C @HailKK

OpenStudy (theedwardsfamily):

i was thinking BC only

OpenStudy (anonymous):

B = 18 is 2/3 of 27, and 20 is 2/3 of 30 C = 27*2, 30*2 D = 27*3, 30*3

OpenStudy (anonymous):

B C and D are all correct because they all include, what you could call multiples of 27/30 b is basically 27/30 divided by 3/3 which is 9/10 which is still the same as 27/30 c is 27/30 times 2/2 d 27/30 times 3/3

OpenStudy (anonymous):

I already explained it...

OpenStudy (anonymous):

So @HailKK you're right

OpenStudy (theedwardsfamily):

ohhhhhhhh. i understand now

OpenStudy (anonymous):

Thanks you!

OpenStudy (holly00d1248):

I agree

OpenStudy (anonymous):

Looks like most of us learned something ;P

OpenStudy (holly00d1248):

LOL

OpenStudy (theedwardsfamily):

medal 4 medal guys? ^_^

OpenStudy (holly00d1248):

what do you mean

OpenStudy (theedwardsfamily):

i give u a medal, u give me a medal

OpenStudy (theedwardsfamily):

cool?

OpenStudy (anonymous):

I did that with a different question yesterday...

OpenStudy (theedwardsfamily):

oh. so no medal? :'(

OpenStudy (anonymous):

I already gave someone a medal

OpenStudy (theedwardsfamily):

lol. ok, no wrries

OpenStudy (theedwardsfamily):

:)

OpenStudy (theedwardsfamily):

@Holly00d1248 ? medal?

OpenStudy (holly00d1248):

I alreay gave it to @HailKK

OpenStudy (anonymous):

So who's still open?

OpenStudy (theedwardsfamily):

Aww man! @Legends ?

OpenStudy (holly00d1248):

what do you mean

OpenStudy (holly00d1248):

@HailKK

OpenStudy (anonymous):

I meant "who still has the privilege to give out a medal?"

OpenStudy (holly00d1248):

i dont know

OpenStudy (theedwardsfamily):

Anybody?????